an 8-foot tall cylindrical water tank has a base of diameter 6 feet.
how much water is in the tank if the water is 5 feet deep?
what is the formula for the volume of water as a function of its depth in the tank?

Answer by richwmiller(9144)

Depth and height are the same thing. Height is looking from the bottom and depth from the top
V=pi*r^2*h
V==pi*3^2*5 volume of water at 5 ft deep.
V=pi*9*5
V=45*pi
pi=about 3.14
radius =1/2 diameter
Why didn't we use 8 ft for the height ?
We didn't fill the tank all the way but only to 5 ft.